PROBLEM 2. A block of wood is in the form of a right circular cone. The altitude is 12 cm and the radius of the base is 5 cm. A cylindrical hole of 5 cm is bored completely through the solid, the axis of the hole coinciding with the axis of the cone. Find the amount of wood left after the hole is bored.
FINAL ANSWER: _______________________________________
1
Expert's answer
2020-12-16T08:17:36-0500

Radius of cone is 5 cm

Height of cone is 12 cm

Radius of cylinder is 2.5 cm

Let height of cylinder is h cm

 By similarity

(12-h)/12 = 2.5/5

=> 12 - h = 6

=> h = 6



Now volume of cone = "\frac{1}{3}π(5²)(12)= 100π cm³"


Volume of cylinder inside = "\u03c0(2.5\u00b2)(6)= 37.5\u03c0 cm\u00b3"

So volume of remaining wood

= (100π - 37.5π ) cm³

= 62.5π cm³

= 196.35 cm³